OpenStudy (kitkat1):

Solve and graph the absolute value inequality: |4x + 1| ≤ 5

OpenStudy (kitkat1):

Solve and graph the absolute value inequality: |4x + 1| ≤ 5. (1 point) number line with closed dots on negative 1.5 and 1 with shading going in the opposite directions. number line with open dots on negative 1.5 and 1 with shading in between. number line with closed dots on negative 1 and 1 with shading in between. number line with closed dots on negative 1.5 and 1 with shading in between.

OpenStudy (kitkat1):

I will medal Please help me!\

OpenStudy (kitkat1):

whpalmer4 can u help me?

OpenStudy (kitkat1):

Can someone plz help me

OpenStudy (kitkat1):

I think its either the 1st one or the last one

OpenStudy (zpupster):

\[-5\le 4x+1\le5\] subtract 1 from all \[-6\le 4x \le 4\]

OpenStudy (zpupster):

divide all by 4 \[-3/2\le x \le 1\]

OpenStudy (kitkat1):

So the answer should be 1.5 and 1

OpenStudy (kitkat1):

Its either the first answer or the second one

OpenStudy (kitkat1):

first one or the last one nevermind